Please show how you got your answer so I can understand what you did:) thanks!Calculate the molar solubility of calcium carbonat
Answer #1

Molarity CaCO3(s)   \rightleftharpoons Ca2+(aq) + CO32-(aq)
Initial 0 0
Change x x
Final x x

Now, Ksp = [Ca2+] [CO32-]

\Rightarrow 4.96 x 10-9 = (x)(x)

\Rightarrow x2 = 4.96 x 10-9

\Rightarrow x = \sqrt{}(4.96 x 10-9)

\Rightarrow x = 7.04 x 10-5

So, molar solubility of CaCO3 = 7.04 x 10-5

Answer is (a)